An airplane in a holding pattern flies at constant altitude along a circular path of radius 3.38 km. If the airplane rounds half the circle in 148 s, determine the following. 


(a) Determine the magnitude of the airplane's displacement during the given time (in m). 

(b) Determine the magnitude of the airplane's average velocity during the given time (in m/s) 

(c) What is the airplane's average speed during the same time interval (in m/s)? 
Answer #1

given the airplane is flying at a constant altitude

radius of circular path, r = 3380 m

time taken to complete one half of a circle, T = 148 s

a. in the given time, the airplane covers half the circel to displacement

d = 2*r = 6760 m

b. average velocity during the given time = d/T

Vav = 6760/148 = 45.675 m/s

c. average speed = pi*r/T

v = pi*3380/148 = 71.747 m/s